# The Impact of Domain Knowledge and Multi-Modality on Intelligent Molecular Property Prediction: A Systematic Survey

### v1

@@ Abstract
The precise prediction of molecular properties is essential for advancements in drug development, particularly in virtual screening and compound optimization. The recent introduction of numerous deep learning-based methods has shown remarkable potential in enhancing molecular property prediction (MPP), especially improving accuracy and insights into molecular structures. Yet, two critical questions arise: does the integration of domain knowledge augment the accuracy of molecular property prediction and does employing multi-modal data fusion yield more precise results than unique data source methods? To explore these matters, we comprehensively review and quantitatively analyze recent deep learning methods based on various benchmarks. We discover that integrating molecular information will improve both MPP regression and classification tasks by upto 3.98% and 1.72%, respectively. We also discover that the utilizing 3-dimensional information with 1-dimensional and 2-dimensional information simultaneously can substantially enhance MPP upto 4.2%. The two consolidated insights offer crucial guidance for future advancements in drug discovery.
@@ 1 Introduction
Despite the advancements in DL for MPP, the field continues to face ongoing evolution and challenges. Two significant trends are currently shaping the field. The first trend is the increasing integration of domain knowledge into DL models. This includes a broad spectrum of scientific information, such as chemical and physical property relation, atom and bond characteristics, and detailed insights into functional groups and molecular fragments.
@@
The integration of this knowledge aims to enhance the predictive accuracy of these models. This leads us to the critical question: Does more comprehensive domain knowledge actually improve the effectiveness of MPP? The second trend is the rising adoption of multi-modality techniques, which involve the fusion of various data types like sequence-based, graph-based and pixel-based formats. This approach is driven by the goal of achieving more accurate predictions in a field characterized by its complexity and data diversity, prompting the question: Is multi-modality more effective for MPP than methodologies that rely on uni-model data source? To explore these questions, our paper begins with an in-depth review of the current DL approaches in MPP, focusing on how the domain knowledge and multi-modal data integrate on encoder architecture and training strategy.
Our review begins with an examination of various unique data encoder architectures for MPP such as Recurrent Neural Network (RNN)[3, 4, 5], Graph Neural Network (GNN)[6, 7, 8, 9, 10], Transformer[11, 12, 13, 14], and Convolutional Neural Network (CNN)[15, 16, 17] models, and also review the multi-modal methods[18, 19, 20, 21] We focus on how these architectures are aligned with existing molecular structural knowledge and their integration of domain knowledge. This exploration highlights the synergy between advanced computational techniques and fundamental molecular understanding, a crucial aspect in enhancing the accuracy of MPP. Also we review a variety of training strategies, such as self-supervised[22, 23, 24], semi-supervised[25, 26, 27], transfer learning[28, 29, 30] and multi-tasks learning[31, 32]. A particular emphasis is placed on strategies that effectively utilize unlabeled data, a vital consideration given the frequent scarcity of labeled data in this domain, and we focus on how the domain knowledge and multi-modal data to be used in the training strategy. Accompanying this review are comprehensive diagrams that systematically elucidate the nuances of these encoder architectures and training strategies, offering a clearer understanding of their complex mechanisms. The overview of our paper is as Figure 1.
@@
$\bullet$ We discover that integrating molecular substructure information results in a 3.98% improvement on average in regression tasks and a 1.72% increase on average in classification tasks.
$\bullet$ We discover that the utilizing 3-dimensional information with 1-dimensional and 2-dimensional information simultaneously can substantially enhance MPP upto 4.2%.
2 Molecular Data
In the field of molecular science, an vast range of molecular data has been developed, each crucial for computational modeling and analysis. These formats are generally classified into three main types: text-based, graph-based, and pixel-based formats. Each type offers unique insights into molecular structures, contributing significantly to various aspects of molecular analysis. These diverse formats are illustrated in the Figure 2, which showcases the array of molecular data available for DL methods.
@@ 2.3 Pixel-based Data
Pixel-Based Molecular Data Formats, such as 2D images and 3D grids, are essential components of molecular property prediction. These formats, easily generated by tools like RDKit[42] and PyMol[43] for 2D images and Libmolgrid[44] for 3D grids, offer clear and comprehensible visual representations of molecular structures. This visual aspect allows for straightforward human interpretation, aiding in the recognition of molecular patterns and the understanding of spatial relationships in computational modeling. This clarity in visualization is crucial for effectively analyzing molecular geometries and interactions.
Figure 2: Molecular data: We illustrate the transformation of molecular data formats crucial for MPP using the example of the caffeine molecule. This is demonstrated across three primary categories: sequence-based, graph-based, and pixel-based formats. Each format is derived from the SMILES representation of caffeine, using Python packages such as RDKit and software tools like PyMol. a). The sequence-based data section includes formats like SMILES, its variants (Canonical and Isomeric SMILES), molecular fingerprints (ECFP, Morgan, MACCS), and SELFIES, highlighting their roles in encoding molecular structures. b). Graph-based data represents caffeine as a graph with atoms as nodes and bonds as edges, enriched with 3D information for detailed structural insights. c). Pixel-based data showcases 2D images and 3D grids of caffeine, crucial for visual analysis and spatial interpretation.
@@ 3 Domain Knowledge
In molecular science, many domain knowledge from areas like physics, chemistry, and biology plays a vital role. This knowledge is methodically grouped into four key categories: atom-bond property, molecular substructure, chemical reactions, and molecular characteristics. Each category is integral for a comprehensive understanding and accurate interpretation of molecular data. The Figure 3 showcases these categories in detail, providing an in-depth look at the essential aspects of molecular information interpretation.
@@ 3.1 Atom-bond Property
In MPP, a deep understanding of atomic and bonding attributes is vital for accurately modeling molecular behaviors. Understanding atomic properties such as isotope numbers, which affect molecular weight and stability, and chirality, crucial for interaction and reaction in biological systems, is essential. Hybridization types impact bonding patterns and molecular geometry. Atomic valence, number, mass, formal charge, and aromaticity significantly influence a molecule’s chemistry.
Figure 3: Molecular domain knowledge: This figure categorizes molecular expert knowledge essential for MPP into four domains, using the molecule C=(CC(=C)c1cc(C(=O)O)cnc1C(C)$C$ as an example. a). In the atom-bond property section, we examine aspects such as the molecule’s atomic number, mass, valence, and bond types, focusing on isotope numbers that influence molecular weight and stability, and chirality affecting how the molecule interacts in biological systems. b). Molecular substructure includes the functional groups, molecular fragments, and pharmacophores of this molecule, illustrating their influence on its chemical behavior and interactions. c). Molecular property covers a range of properties from quantum mechanics to physiology, showcasing how these properties affect the molecule’s behavior in drug development. d). Chemical reaction discusses the mechanisms of molecular transformations, highlighting the molecule’s reactivity.
Bond attributes like bond type and stereochemistry are key in determining molecular connectivity, reactivity, and shape, influencing interactions with biological targets. The direction and length of bonds also provide insights into spatial arrangement. These detailed atomic and bond attributes collectively provide a comprehensive framework for molecular structure analysis, essential for effective predictive modeling in drug discovery.
@@ 3.2 Molecular Substructure
A pharmacophore is an abstract representation of the molecular features necessary for a molecule to interact with a specific biological target to produce a desired biological effect. A pharmacophore with both a hydrogen bond donor and an acceptor in a specific spatial arrangement, for instance, can be crucial for binding to biological targets like enzymes or receptors, influencing the molecule’s effectiveness as a therapeutic agent. The accurate identification and understanding of these substructures are key to developing new pharmaceuticals, offering detailed insights into molecular interactions.
@@ 3.4 Molecular Property
MPP in drug discovery is a multidisciplinary field, each discipline offering detailed insights into molecular behavior. Quantum mechanics, for example, delves into electronic properties like ionization potentials, crucial for understanding reaction mechanisms. Physical chemistry examines molecules’ stability, reactivity, and phase behaviors, impacting drug formulation. Biophysics explores molecular interactions within biological systems, crucial for drug-target binding studies. Physiology, on the other hand, assesses drug effects at an organismal level, influencing pharmacodynamics and pharmacokinetics. These interconnected properties, such as how a drug’s solubility impacts absorption and bioavailability, highlight the need for a comprehensive understanding across levels, from atomic to organismal, to predict molecular properties accurately and develop effective pharmaceuticals. This integrative approach, encompassing everything from electron distribution to organismal response, is vital in the nuanced field of drug development.
@@ 4.1 Encoder
In MPP, encoder architectures play a key role in transforming raw molecular data into meaningful representations. This section examines a variety of encoder architectures, each tailored to specific molecular data formats and complexities. We categorize four main types of encoders for single data sources: RNN-based, GNN-based, Transformer-based, and CNN-based. Each type is analyzed for its alignment with molecular prior structural knowledge and the integration of domain-specific information. Additionally, we examine multi-modality based encoders, which handle multiple data sources, highlighting their unique characteristics, applications, and the challenges they address in molecular representation learning. The detailed aspects of these encoder architectures are illustrated in the Figure 4.
@@ 4.1.1 RNN-based
Figure 4: Molecular encoder architectures: This figure categorizes molecular encoder architectures into four types: RNN-based, Transformer-based, GNN-based, and CNN-based. Each type is assessed for its ability to capture information about functional groups like carboxyl groups (-COOH) in the molecule C=(CC(=C)c1cc(C(=O)O)cnc1C(C)$C$. a). RNN-based encoders process sequence data, maintaining a memory of previous inputs to effectively capture sequential patterns of (-COOH). b). Transformer-based models utilize Self-Attention mechanisms, enabling them to identify and focus on the (-COOH) group’s specific interactions within the molecular sequence. c). GNN-based architectures employ a message passing strategy, extracting the topological information of (-COOH) within the molecule’s graph structure. d). CNN-based models analyze spatial patterns through convolution layers, identifying sub-images that contain the (-COOH) group. This visualization highlights how each encoder type uniquely processes and interprets the molecular structure for MPP.
RNN, like long short-term memory (LSTM)[50] and gated recurrent unit (GRU)[51], are adept at processing sequential data, with a unique internal memory feature that allows them to maintain context and order in sequences. This capability makes RNNs highly effective for tasks involving sequences data. Nowaday some work use RNN-based model to analyze 1D molecular data such as SMILES.
@@ 4.1.2 GNN-based
But these methods focus on atom (node) or bond (edge) information. To address this issue, Song et al.[8] propose a Communicative Message Passing Neural Network to improve molecular embedding by strengthening the message interactions between nodes and edges through a communicative kernel. SC-NMP[59] aggregates the node representations of the current step and the graph representation of the previous step, and proposes densely self-connected neural message passing, which connects each layer to every other layer in a feed-forward fashion. To extract useful interactions between a target atom and its neighboring atomic groups, Li et al.[60] propose a new graph learning paradigm based on a block design named block-based GNN and demonstrate that the network degradation problem can be reduced by applying block design with normalization and skip-connection. Ma et al.[61] employ cross-dependent message passing strategy to integrate the node-centered and edge-centered encoders. Liu et al.[62] develop a hypergraph-based topological framework to characterize detailed molecular structures and interactions at the atomic level. They have recently proposed embedding homology and persistent homology. Feng et al.[63] transform each molecular graph into a heterogeneous atom-bond graph to fully utilize the bond attributes and design unidirectional position encoding for such graphs. Biswas[31] pass additional atomic and molecular features, including 2D RDKit descriptors, Abraham parameters, QM descriptors, and 3D geometries, to improve the model performance.
{forest}
Figure 5: The molecular encoder architecture summary.
Hasebe[71] proposed a knowledge-embedded message passing nerual network that can be supervised together with nonquantitative knowledge annotations by human experts on a chemical graph. This graph contains information on the important substructure of a molecule and its effect on the target property. Yang et al.[72] extract physical information with a neural physical engine that learns molecular conformations by simulating molecular dynamics with parameterized forces. They then employ this physical information as supplementary data for predicting molecular properties.
@@
However, most methods essentially attribute predictions to individual nodes, edges, or node features. This kind of interpretability is only partially compatible with chemists’ intuition at best. Chemists are more accustomed to comprehending the causal relationship between molecular structures and properties in terms of chemically meaningful substructures, such as functional groups, rather than individual atoms or bonds. Zang et al.[73] decompose the molecular graph by BRICS and additional decomposition to construct a motif-level graph, in which corresponding multi-level generative and predictive tasks are designed as self-supervised signals. As the graph pooling technique for learning expressive graph-level representation is critical yet still challenging, Liu et al.[74] propose master-orthogonal attention, a novel cross-level attention mechanism specifically designed for hierarchical graph pooling. To fully explore higher-order substructure information, Gao et al.[75] propose substructure interaction attention, which takes both the information of neighbors’ substructures and the interaction information among them into account during the aggregation process. To retain locality and linear network complexity, Bouritsas et al.[7] employ a topologically-aware message passing scheme based on substructure encoding, which does not attempt to adhere to the Weisfeiler-Leman hierarchy. Addressing the oversmoothing problem in multi-hop operations, Ye et al.[76] construct a composite molecular representation with multi-substructural feature extraction and process such features effectively with a nested convolution plus readout scheme to capture interacting substructural information. Zhu[77] utilize corepresentation learning of molecular graphs and chemically synthesizable BRICS fragments. Furthermore, a plug-and-play feature-wise attention block is first designed in the their model architecture to adaptively recalibrate atomic features after the message passing phase. To accurately model the complex quantum interactions inherent in molecules, Lu et al.[78] utilize a sophisticated hierarchical graph neural network, which directly extracts features from both the conformation and spatial information of molecules, and then integrates these features through multilevel interactions. Fey et al.[79] take in two complementary graph representations: the raw molecular graph representation and its associated junction tree, where nodes represent meaningful clusters in the original graph. Focusing on the molecular hierarchical relationship, Han et al.[6] propose a simple yet effective rescaling module, called contextual self-rescaling, that adaptively recalibrates molecular representations by explicitly modeling interdependencies between atom and motif features. Ji et al.[64] model a molecule as a heterogeneous graph and leverage metapaths to capture latent features for chemical functional groups. They also design a hierarchical attention strategy to aggregate heterogeneous information at both the node and relation levels. To extract functional groups as motifs for small molecules, Wu et al.[80] construct a heterogeneous molecular graph with both atom-level and motif-level nodes and adopt a heterogeneous self-attention layer to distinguish the interactions between multi-level nodes.
@@
Since different 3D structures may lead to dissimilar molecular properties despite having the same 2D molecular topology, Recently, many works utilizing molecular 3D structures have been introduced. To emphasize equivariant constraints, Fuchs et al.[81] utilize the explicit increase of equivariance constraints in self-attention mechanisms. As rotation-invariant representations struggle to convey directional information, Schutt et al.[82] proposed rotationally equivariant message passing, exemplified by the Polarizable Atom Interaction Neural Network architecture. Brandstetter et al.[83] expand equivariant graph networks to include not only invariant scalar attributes but also covariant information like vectors or tensors. This model consists of steerable MLPs, capable of incorporating geometric and physical information within its message and update functions. Gasteiger et al.[84] show the universality of spherical representations and employ a two-hop message passing mechanism with directed edge embeddings for rotationally equivariant predictions, and utilize symmetric message passing, augmented with geometric information, to enhance our model’s efficacy in MPP. Gasteiger et al.[85] integrate directional information and interatomic distances by embedding and updating messages between atoms, using a spherical 2D Fourier-Bessel basis to jointly represent distances and angles. To model angular relationships among neighboring atoms in a GNN, ensuring constraints like rotation invariance and energy conservation, Shuaibi et al.[86] utilize a per-edge local coordinate frame and innovate a spin convolution, thereby securing rotation invariance in edge messaging. Fang et al.[65] proposed a self-supervised framework using molecular geometric information by constructing a new bond angle graph, where the chemical bonds within a molecule are considered as nodes and the angle formed between two bonds is considered as the edge.

### v3

# Impact of Domain Knowledge and Multi-Modality on Intelligent Molecular Property Prediction: A Systematic Survey

@@ Abstract
The precise prediction of molecular properties is essential for advancements in drug development, particularly in virtual screening and compound optimization. The recent introduction of numerous deep learning-based methods has shown remarkable potential in enhancing molecular property prediction (MPP), especially improving accuracy and insights into molecular structures. Yet, two critical questions arise: does the integration of domain knowledge augment the accuracy of molecular property prediction and does employing multi-modal data fusion yield more precise results than unique data source methods? To explore these matters, we comprehensively review and quantitatively analyze recent deep learning methods based on various benchmarks. We discover that integrating molecular information significantly improves molecular property prediction (MPP) for both regression and classification tasks. Specifically, regression improvements, measured by reductions in root mean square error (RMSE), are up to 4.0%, while classification enhancements, measured by the area under the receiver operating characteristic curve (ROC-AUC), are up to 1.7%. We also discover that enriching 2D graphs with 1D SMILES boosts multi-modal learning performance for regression tasks by up to 9.1%, and augmenting 2D graphs with 3D information increases performance for classification tasks by up to 13.2%, with both enhancements measured using ROC-AUC. The two consolidated insights offer crucial guidance for future advancements in drug discovery.
@@ 1 Introduction
Despite the advancements in DL for MPP, the field continues to face ongoing evolution and challenges. Two significant trends are currently shaping the field. The first trend is the increasing integration of domain knowledge into DL models. This includes a broad spectrum of scientific information, such as chemical and physical property relation, atom and bond characteristics, and detailed insights into functional groups and molecular fragments. The integration of this knowledge aims to enhance the predictive accuracy of these models. This leads us to the critical question: Does more comprehensive domain knowledge actually improve the effectiveness of MPP? The second trend is the rising adoption of multi-modality techniques, which involve the fusion of various data types like sequence-based, graph-based and pixel-based formats. This approach is driven by the goal of achieving more accurate predictions in a field characterized by its complexity and data diversity, prompting the question: Is multi-modality more effective for MPP than methodologies that rely on uni-model data source? To explore these questions, our paper begins with an in-depth review of the current DL approaches in MPP, focusing on how the domain knowledge and multi-modal data integrate on encoder architecture and training strategy.
Our review begins with an examination of various unique data encoder architectures for MPP such as Recurrent Neural Network (RNN)[3, 4, 5], Graph Neural Network (GNN)[6, 7, 8, 9, 10], Transformer[11, 12, 13, 14], and Convolutional Neural Network (CNN)[15, 16, 17] models, and also review the multi-modal methods[18, 19, 20, 21]. We focus on how these architectures are aligned with existing molecular structural knowledge and their integration of domain knowledge. This exploration highlights the synergy between advanced computational techniques and fundamental molecular understanding, a crucial aspect in enhancing the accuracy of MPP. Also we review a variety of training strategies, such as self-supervised[22, 23, 24], semi-supervised[25, 26, 27], transfer learning[28, 29, 30] and multi-tasks learning[31, 32]. A particular emphasis is placed on strategies that effectively utilize unlabeled data, a vital consideration given the frequent scarcity of labeled data in this domain, and we focus on how the domain knowledge and multi-modal data to be used in the training strategy. Accompanying this review are comprehensive diagrams that systematically elucidate the nuances of these encoder architectures and training strategies, offering a clearer understanding of their complex mechanisms. The overview of our paper is as Figure 1.
@@
$\bullet$ We discover that integrating molecular substructure information results in a 4.0% improvement on average in regression tasks and a 1.7% increase on average in classification tasks.
$\bullet$ We discover that enriching 2D graph models with 1D SMILES or 3D information boosts multi-modal learning, enhancing performance by 9.1% to 13.2% over single-modality models.
2 Molecular Modality
In the field of molecular science, an vast range of molecular modality has been developed, each crucial for computational modeling and analysis. These formats are generally classified into three main types: text-based, graph-based, and pixel-based formats. Each type offers unique insights into molecular structures, contributing significantly to various aspects of molecular analysis. These diverse formats are illustrated in the Figure 2, which showcases the array of molecular modality available for DL methods.
@@ 2.3 Pixel-based Data
Pixel-Based Molecular Data Formats, such as 2D images and 3D grids, are essential components of molecular property prediction. These formats, easily generated by tools like RDKit[42] and PyMol[43] for 2D images and Libmolgrid[44] for 3D grids, offer clear and comprehensible visual representations of molecular structures. This visual aspect allows for straightforward human interpretation, aiding in the recognition of molecular patterns and the understanding of spatial relationships in computational modeling.
Figure 2: Molecular modality: We illustrate the transformation of molecular modality crucial for MPP using the example of the caffeine molecule. This is demonstrated across three primary categories: sequence-based, graph-based, and pixel-based formats. Each format is derived from the SMILES representation of caffeine, using Python packages such as RDKit and software tools like PyMol. a). The sequence-based data section includes formats like SMILES and its variants (Canonical and Isomeric SMILES), molecular fingerprints (ECFP, Morgan, MACCS), and SELFIES, highlighting their roles in encoding molecular structures. b). Graph-based data represents caffeine as a graph with atoms as nodes and bonds as edges, enriched with 3D information for detailed structural insights. c). Pixel-based data showcases 2D images and 3D grids of caffeine, crucial for visual analysis and spatial interpretation.
This clarity in visualization is crucial for effectively analyzing molecular geometries and interactions.
@@ 3 Domain Knowledge
In molecular science, many domain knowledge from areas, like physics, chemistry, and biology, play a vital role. This knowledge is methodically grouped into four key categories: atom-bond property, molecular substructure, chemical reactions, and molecular characteristics. Each category is integral for a comprehensive understanding and accurate interpretation of molecular data. The Figure 3 showcases these categories in detail, providing an in-depth look at the essential aspects of molecular information interpretation.
@@ 3.1 Atom-bond Property
In MPP, a deep understanding of atomic and bonding attributes is vital for accurately modeling molecular behaviors. Understanding atomic properties is essential for molecular analysis. For example, isotope numbers influence molecular weight and stability, and chirality is crucial for interactions and reactions within biological systems. Hybridization types impact bonding patterns and molecular geometry.
Figure 3: Molecular domain knowledge: This figure categorizes molecular expert knowledge essential for MPP into four domains, using the molecule C=(CC(=C)c1cc(C(=O)O)cnc1C(C)$C$ as an example. a). In the atom-bond property section, we examine aspects such as the molecule’s atomic number, mass, valence, and bond types. b). Molecular substructure includes the functional groups, molecular fragments, and pharmacophores of this molecule, illustrating their influence on its chemical behavior and interactions. c). Molecular property covers a range of properties from quantum mechanics to physiology, showcasing how these properties affect the molecule’s behavior in drug development. d). Chemical reaction discusses the mechanisms of molecular transformations, highlighting the molecule’s reactivity.
Atomic valence, number, mass, formal charge, and aromaticity significantly influence a molecule’s chemistry. Bond attributes like bond type and stereochemistry are key in determining molecular connectivity, reactivity, and shape, influencing interactions with biological targets. The direction and length of bonds also provide insights into spatial arrangement. These detailed atomic and bond attributes collectively provide a comprehensive framework for molecular structure analysis, essential for effective predictive modeling in drug discovery.
@@ 3.2 Molecular Substructure
A pharmacophore is an abstract representation of the molecular features that are necessary for a molecule to interact with a specific biological target to produce a desired biological effect. A pharmacophore with both a hydrogen bond donor and an acceptor in a specific spatial arrangement, for instance, can be crucial for binding to biological targets like enzymes or receptors, influencing the molecule’s effectiveness as a therapeutic agent. The accurate identification and understanding of these substructures are key to developing new pharmaceuticals, offering detailed insights into molecular interactions.
@@ 3.4 Molecular Property
MPP in drug discovery is a multidisciplinary field, each discipline offering detailed insights into molecular behavior. Quantum mechanics, for example, delves into electronic properties like ionization potentials, crucial for understanding reaction mechanisms. Physical chemistry examines the stability of molecule, reactivity, and phase behaviors, impacting drug formulation. Biophysics explores molecular interactions within biological systems, crucial for drug-target binding studies. Physiology, on the other hand, assesses drug effects at an organismal level, influencing pharmacodynamics and pharmacokinetics. These interconnected properties, such as how a drug’s solubility impacts absorption and bioavailability, highlight the need for a comprehensive understanding across levels, from atomic to organismal, to predict molecular properties accurately and develop effective pharmaceuticals. This integrative approach, encompassing everything from electron distribution to organismal response, is vital in the nuanced field of drug development.
@@ 4.1 Encoder
In MPP, encoder architectures play a key role in transforming raw molecular data into meaningful representations. This section examines a variety of encoder architectures, each tailored to specific molecular modality and complexities. We categorize four main types of encoders for single data sources: RNN-based, GNN-based, Transformer-based, and CNN-based. Each type is analyzed for its alignment with molecular prior structural knowledge and the integration of domain-specific information. Additionally, we examine multi-modality based encoders, which handle multiple data sources, highlighting their unique characteristics, applications, and the challenges they address in molecular representation learning. The detailed aspects of these encoder architectures are illustrated in the Figure 4 and Figure 5.
{forest}
Figure 4: The molecular encoder method summary. We categorize molecular encoder method into five types: RNN-based, GNN-based, Transformer-based, CNN-based, and Multi-Modality-based. For each category, key techniques and notable advancements utilized in various influential studies are highlighted, showcasing the evolution and diversification of approaches in molecular encoding.
@@ 4.1.1 RNN-based
RNN, like long short-term memory (LSTM)[59] and gated recurrent unit (GRU)[60], are adept at processing sequential data, with a unique internal memory feature that allows them to maintain context and order in sequences. This capability makes RNNs highly effective for tasks involving sequences data. Nowadays some work uses RNN-based model to analyze 1D molecular data, such as SMILES.
@@ 4.1.2 GNN-based
But these methods focus on atom (node) or bond (edge) information. To address this issue, Song et al.[8] propose a Communicative Message Passing Neural Network to improve molecular embedding by strengthening the message interactions between nodes and edges through a communicative kernel. SC-NMP[66] aggregates the node representations of the current step and the graph representation of the previous step, and proposes densely self-connected neural message passing, which connects each layer to every other layer in a feed-forward fashion. To extract useful interactions between a target atom and its neighboring atomic groups, Li et al.[67] proposed a new graph learning paradigm based on a block design named block-based GNN and demonstrated that the network degradation problem can be reduced by applying a block design with normalization and skip-connection. Ma et al.[68] employ cross-dependent message passing strategy to integrate the node-centered and edge-centered encoders. Liu et al.[69] develop a hypergraph-based topological framework to characterize detailed molecular structures and interactions at the atomic level. They have recently proposed embedding homology and persistent homology. Feng et al.[70] transform each molecular graph into a heterogeneous atom-bond graph to fully utilize the bond attributes and design unidirectional position encoding for such graphs. Biswas[31] pass additional atomic and molecular features, including 2D RDKit descriptors, Abraham parameters, QM descriptors, and 3D geometries, to improve the model performance. Hasebe[71] proposed a knowledge-embedded message passing nerual network that can be supervised together with nonquantitative knowledge annotations by human experts on a chemical graph. This graph contains information on the important substructure of a molecule and its effect on the target property. Yang et al.[72] extract physical information with a neural physical engine that learns molecular conformations by simulating molecular dynamics with parameterized forces. They then employ this physical information as supplementary data for predicting molecular properties.
@@
However, most methods essentially attribute predictions to individual nodes, edges, or node features. This kind of interpretability is only partially compatible with chemists’ intuition at best. Chemists are more accustomed to comprehending the causal relationship between molecular structures and properties in terms of chemically meaningful substructures, such as functional groups, rather than individual atoms or bonds. Zang et al.[73] decompose the molecular graph by BRICS and additional decomposition to construct a motif-level graph, in which corresponding multi-level generative and predictive tasks are designed as self-supervised signals. As the graph pooling technique for learning expressive graph-level representation is critical yet still challenging, Liu et al.[74] propose master-orthogonal attention, a novel cross-level attention mechanism specifically designed for hierarchical graph pooling. To fully explore higher-order substructure information, Gao et al.[75] propose substructure interaction attention, which takes both the information of neighbors’ substructures and the interaction information among them into account during the aggregation process. To retain locality and linear network complexity, Bouritsas et al.[7] employ a topologically-aware message passing scheme based on substructure encoding, which does not attempt to adhere to the Weisfeiler-Leman hierarchy. Addressing the oversmoothing problem in multi-hop operations, Ye et al.[76] construct a composite molecular representation with multi-substructural feature extraction and process such features effectively with a nested convolution plus readout scheme to capture interacting substructural information. Zhu[77] utilize corepresentation learning of molecular graphs and chemically synthesizable BRICS fragments. Furthermore, a plug-and-play feature-wise attention block is first designed in the their model architecture to adaptively recalibrate atomic features after the message passing phase. To accurately model the complex quantum interactions inherent in molecules, Lu et al.[78] utilize a sophisticated hierarchical graph neural network, which directly extracts features from both the conformation and spatial information of molecules, and then integrates these features through multilevel interactions. Fey et al.[79] take in two complementary graph representations: the raw molecular graph representation and its associated junction tree, where nodes represent meaningful clusters in the original graph.
Figure 5: Molecular encoder architectures: This figure categorizes molecular encoder architectures for single modality into four types: RNN-based, Transformer-based, GNN-based, and CNN-based. Each type is assessed for its ability to capture information about functional groups like carboxyl groups (-COOH) in the molecule C=(CC(=C)c1cc(C(=O)O)cnc1C(C)$C$. a). RNN-based encoders process sequence data, maintaining a memory of previous inputs to effectively capture sequential patterns of (-COOH). b). Transformer-based models utilize Self-Attention mechanisms, enabling them to identify and focus on the (-COOH) group’s specific interactions within the molecular sequence. c). GNN-based architectures employ a message passing strategy, extracting the topological information of (-COOH) within the molecule’s graph structure. d). CNN-based models analyze spatial patterns through convolution layers, identifying sub-images that contain the (-COOH) group. This visualization highlights how each encoder type uniquely processes and interprets the molecular structure for MPP.
Focusing on the molecular hierarchical relationship, Han et al.[6] propose a simple yet effective rescaling module, called contextual self-rescaling, that adaptively recalibrates molecular representations by explicitly modeling interdependencies between atom and motif features. Ji et al.[52] model a molecule as a heterogeneous graph and leverage metapaths to capture latent features for chemical functional groups. They also design a hierarchical attention strategy to aggregate heterogeneous information at both the node and relation levels. To extract functional groups as motifs for small molecules, Wu et al.[80] construct a heterogeneous molecular graph with both atom-level and motif-level nodes and adopt a heterogeneous self-attention layer to distinguish the interactions between multi-level nodes.
@@
Since different 3D structures may lead to dissimilar molecular properties despite having the same 2D molecular topology, Recently, many works utilizing molecular 3D structures have been introduced. To emphasize equivariant constraints, Fuchs et al.[81] utilize the explicit increase of equivariance constraints in self-attention mechanisms. As rotation-invariant representations struggle to convey directional information, Schutt et al.[82] proposed rotationally equivariant message passing, exemplified by the Polarizable Atom Interaction Neural Network architecture. Brandstetter et al.[83] expand equivariant graph networks to include not only invariant scalar attributes but also covariant information like vectors or tensors. This model consists of steerable MLPs, capable of incorporating geometric and physical information within its message passing and update functions. Gasteiger et al.[84] show the universality of spherical representations and employ a two-hop message passing mechanism with directed edge embeddings for rotationally equivariant predictions, and utilize symmetric message passing, augmented with geometric information, to enhance our model’s efficacy in MPP. Gasteiger et al.[85] integrate directional information and interatomic distances by embedding and updating messages between atoms, using a spherical 2D Fourier-Bessel basis to jointly represent distances and angles. To model angular relationships among neighboring atoms in a GNN, ensuring constraints like rotation invariance and energy conservation, Shuaibi et al.[86] utilize a per-edge local coordinate frame and innovate a spin convolution, thereby securing rotation invariance in edge messaging. Fang et al.[53] proposed a self-supervised framework using molecular geometric information by constructing a new bond angle graph, where the chemical bonds within a molecule are considered as nodes and the angle formed between two bonds is considered as the edge.